how do you solve the system by graphing 2x+y=0
6x+4y=6

Answer by Alan3354(69443)

how do you solve the system by graphing 2x+y=0
6x+4y=6
--------------
Both equations are straight lines, so you need 2 points for each.
2x+y = 0
y = -2x
Pick 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.
========================
6x+4y=6
3x+2y=3
y = (-3/2)x + 3/2
Do the same for this one.
------------
Every point on the line is a solution for the eqn.
The intersection is a solution for both, the solution of the system of equations.
-----------
Try it, email me via the Thank You note to check your work, or for help.